Find the sum of n terms :-
12+42+72+.............
Question based on Sequence and series.

Grade:11

2 Answers

Arun
6 years ago
its n terms will be (2n –1)2
thus Tn = 4n^2 + 1 – 4n
so Sn = \sumTn
Sn  = \sum(4n^2 + 1 – 4n)
now open sigma seperately
Sn = \sum 4n^2 + \sum 1 – \sum 4n
 
put formula and you will get the answer
 
Hope it helps
SHISHIR KUMAR PADHAN
3 years ago
Its n term will be (3n-2)
So the terms can be written as,
9n2+12n+4,
S S= 9(12+22+32+...+n)+12(1+2+3+...+n)+4
= {9(n)(n+1)(2n+1)/6}+{12(n)(n+1)/2}+4